Cryptology Assignment 2

Due Monday, March 25, 1996.

• Do problems 4.11, 4.14, and 4.15 in the textbook. (For 4.14, skip part (b), but assume the result there for the later subquestions.) In problem 4.11, it should say “having success probability”, rather than “having failure probability”. In 4.15c, the floor symbol should be a ceiling.

• Calculate the Jacobi symbols (\frac{16}{57}), (\frac{20}{59}), and (\frac{23}{59}).